### 10 Apr 2024 v4

# Zipformer: A faster and better encoder for automatic speech recognition

Abstract
The Conformer has become the most popular encoder model for automatic speech recognition (ASR). It adds convolution modules to a Transformer to learn both local and global dependencies. In this work we describe a faster, more memory-efficient, and better-performing Transformer, called Zipformer. Modeling changes include: 1) a U-Net-like encoder structure where middle stacks operate at lower frame rates; 2) reorganized block structure with more modules, within which we re-use attention weights for efficiency; 3) a modified form of LayerNorm called BiasNorm allows us to retain some length information; 4) new activation functions SwooshR and SwooshL work better than Swish. We also propose a new optimizer, called ScaledAdam, which scales the update by each tensor’s current scale to keep the relative change about the same, and also explictly learns the parameter scale. It achieves faster convergence and better performance than Adam. Extensive experiments on LibriSpeech, Aishell-1, and WenetSpeech datasets demonstrate the effectiveness of our proposed Zipformer over other state-of-the-art ASR models. Our code is publicly available at https://github.com/k2-fsa/icefall.
1 Introduction
End-to-end models have achieved remarkable success in automatic speech recognition (ASR). An effective encoder architecture that performs temporal modeling on the speech sequence plays a vital role in end-to-end ASR models. A most prominent example is Conformer (Gulati et al., 2020), which combines the advantages of the convolutional neural network (CNN) models (Zhang et al., 2017; Li et al., 2019; Kriman et al., 2020) and Transformer models (Dong et al., 2018; Karita et al., 2019; Zhang et al., 2020b). By integrating CNN into Transformer (Vaswani et al., 2017), Conformer is able to extract both local and global dependencies on speech sequences, and achieves state-of-the-art performance in ASR.
In this work, we propose a faster, more memory-efficient, and better-performing Transformer as ASR encoder, called Zipformer. First, unlike Conformer that operates on the sequence at a constant frame rate, Zipformer adopts a U-Net-like (Ronneberger et al., 2015) structure, which consists of multiple stacks downsamping the sequence to various lower frame rates. Second, we re-design the block structure, which is equipped with more modules like two Conformer blocks, and reuses the attention weights for efficiency. We propose BiasNorm as a simpler replacement of LayerNorm, which allows for retaining length information in normalization. We also replace Swish with our new activation functions SwooshR and SwooshL to achieve better results. In addition, we devise a parameter-scale-invariant version of Adam, called ScaledAdam, which scales the update by the current parameter scale and also explicitly learns the parameter scale. Compared to Adam, ScaledAdam enables faster convergence and better performance.
Extensive experiments are conducted on LibriSpeech, Aishell-1, and WenetSpeech datasets, and results demonstrate the effectiveness of the proposed modeling and optimization-related innovations. Zipformer achieves state-of-the-art results on all three datasets. It is worth mentioning that Zipformer is the first model ever to achieve results comparable to those reported in the Conformer paper on the LibriSpeech dataset (these results have proved difficult for others to reproduce). In terms of efficiency, Zipformer converges faster during training and speeds up the inference by more than 50% compared to previous studies while requiring less GPU memory. We perform detailed ablation studies to investigate the contribution of individual components.
2 Related Work
Model architecture. Deep convolution architectures have been applied to end-to-end ASR (Zhang et al., 2017; Li et al., 2019). Follow-up works explore improvements by using depthwise separable convolutions (Howard et al., 2017) for efficiency (Kriman et al., 2020), and incorporating squeeze-and-excitation module (Hu et al., 2018) to capture longer context (Han et al., 2020). Inspired by the success of Transformer (Vaswani et al., 2017) in natural language processing (NLP) field, some works adapt Transformer to speech applications (Dong et al., 2018; Karita et al., 2019; Zhang et al., 2020b; Wang et al., 2020; Zhang et al., 2020a). Compared to CNN, the remarkable benefit of Transformer is that it can learn global dependencies based on self-attention, which is essential for speech processing task. By integrating convolution into Transformer, Conformer (Gulati et al., 2020) gains powerful capability of modeling both local and global contextual information, and outperforms all previous ASR models.
Recent works explore architecture changes on Conformer to further reduce the computational cost and improve the recognition performance. Squeezeformer (Kim et al., 2022) adopts a temporal U-Net structure in which the middle modules operate at half frame rates, and also redesigns the block structure to make it similar to the standard Transformer block (Vaswani et al., 2017). Branchformer (Peng et al., 2022) incorporates parallel branches to model various ranged context, in which one branch captures the local context with convolutional gating multi-layer perceptron (MLP), while the other branch learns long-range dependencies with self-attention. E-Branchformer (Kim et al., 2023) further improves Branchformer by enhancing the branch merging mechanism by convolution-based module.
Zipformer shares similar ideas about temporal downsampling as the previous work Squeezeformer. However, compared to the fixed downsampling ratio in Squeezeformer, Zipformer operates at different downsampling ratios at different encoder stacks and uses much more aggressive downsampling ratios in the middle encoder stacks. In addition to the modeling differences, our work also focuses on optimization-related changes including a new optimizer ScaledAdam, which are shown to improve convergence in the experiments.
End-to-end framework. Connectionist temporal classification (CTC) (Graves et al., 2006) is one of the earliest frameworks for end-to-end ASR, but its performance is limited by the frame independent assumption. To this end, a hybrid architecture that integrates attention-based encoder-deocder (AED)  (Chan et al., 2015) in CTC (Watanabe et al., 2017) (CTC/AED) is proposed to improve the performance. Neural transducer (Graves, 2012), commonly known as RNN-T, addresses the frame independence assumption using a label decoder and a joint network and becomes a popular framework due to its superior performance. Recently, various approaches such as pruning (Kuang et al., 2022; Wang et al., 2023; Mahadeokar et al., 2021) or batch-splitting (Kuchaiev et al., 2019) are proposed to accelerate the training speed and reduce memory usage of neural transducers.
3 Method
3.1 Downsampled encoder structure
Figure 1 presents the overall architecture of the proposed Zipformer model. Different from Conformer (Gulati et al., 2020) that processes the sequence at a fixed frame rate of 25Hz, Zipformer uses a U-Net-like structure learning temporal representation at different resolutions in a more efficient way. Specifically, given the acoustic features with frame rate of 100Hz, the convolution-based module called Conv-Embed first reduces the length by a factor of 2, resulting in a 50Hz embedding sequence. The obtained sequence is then fed into 6 cascaded stacks to learn temporal representation at frame rates of 50Hz, 25Hz, 12.5Hz, 6.25Hz, 12.5Hz, and 25Hz, respectively. Except for the first stack, the other stacks all adopt the downsampled structures, processing the sequence at lower frame rates. The frame rate between stacks is consistently 50Hz. Different stacks have different embedding dimensions, and the middle stacks have larger dimensions. The output of each stack is truncated or padded with zeros to match the dimension of the next stack. The final encoder output dimension is set to the maximum of all stacks’ dimensions. Specifically, if the last stack output has the largest dimension, it is taken as the encoder output; otherwise, it is concatenated from different pieces of stack outputs, taking each dimension from the most recent output that has it present. Finally, a Downsample module converts the sequence to 25Hz, resulting in the encoder output.
Conv-Embed. In Conv-Embed we use three 2-D convolutional layers with time $\times$ frequency strides of $1\times 2$, $2\times 2$, and $1\times 2$, and output channels of 8, 32, and 128, respectively. Subsequently, we utilize one ConvNeXt layer (Liu et al., 2022) similar to Nextformer (Jiang et al., 2022), which is composed of a depth-wise convolution with kernel size of $7\times 7$, a point-wise convolution with 384 output channels, a SwooshL activation function (described in Section 3.4), and a point-wise convolution with 128 output channels. Residual connection is applied on the ConvNeXt module. Finally, a linear layer followed by a BiasNorm (described in Section 3.3) is used to adjust the feature dimension to match the first stack.
Downsampled stacks. In the downsampled stacks, the pairwise Downsample and Upsample modules perform symmetric scaling down and scaling up in sequence length, respectively, using almost the simplest methods. For example, with a factor of 2, the Downsample module averages every 2 frames with 2 learnable scalar weights (after softmax normalization), and the Upsample module just repeats each frame twice. After downsampling, it employs the stacking Zipformer blocks (described in Section 3.2) for temporal modeling at lower frame rates. Finally, it utilizes the Bypass module (described in Section 3.2) to combine the stack input and stack output in a learnable way.
3.2 Zipformer block
Figure 1: Overall architecture of Zipformer.
Figure 2: (Left): Zipformer block structure. (Right): Non-Linear Attention module structure.
Conformer block consists of four modules: feed-forward, Multi-Head Self-Attention (MHSA), convolution, and feed-forward. MHSA learns global context by two steps: computing attention weights using the dot-product operation and aggregating different frames with these attention weights. However, MHSA typically accounts for a large computational cost, since above two steps both require quadratic complexity with respect to the sequence length. Hence, we decompose MHSA into two individual modules according to above two steps: Multi-Head Attention Weight (MHAW) and Self-Attention (SA). This change allows to perform the attention computation twice more efficiently in each block by using one MHAW module and two SA modules. In addition, we propose a new module called Non-Linear Attention (NLA) to make full use of the computed attention weights to capture the global information.
As illustrated in Figure 2 (Left), Zipformer block is equipped with about twice the depth of the Conformer block (Gulati et al., 2020). The main motivation is to allow the re-use of the attention weights to save time and memory. Specifically, the block input is first fed into an MHAW module, which calculates the attention weights and shares them with an NLA module and two SA modules. Meanwhile, the block input is also fed into a feed-forward module followed by the NLA module. Then it applies two module groups, each consisting of SA, convolution, and feed-forward. Finally, a BiasNorm (described in Section 3.3) is used to normalize the block output. In addition to the regular residual connections using adding operation, each block utilizes two Bypass modules to combine the block input and the module outputs, placed in the middle and end of the block. Note that different from regular Transformer models (Vaswani et al., 2017), we don’t use normalization layer such as LayerNorm (Ba et al., 2016) for each module to periodically prevent activations from becoming either too large or too small, since our proposed ScaledAdam optimizer is able to learn the parameter scales (described in Section 3.5).
Non-Linear Attention. Figure 2 (Right) presents the NLA structure. It also leverages the pre-computed attention weights from MHAW to aggregate the embedding vectors over the time axis, which is similar to SA. Specifically, it first projects the input with 3 linear layers to $A$, $B$, and $C$, each being of 3/4 input dimension. The module output is $\mathrm{linear}(A\odot\mathrm{attention}(\tanh(B)\odot C))$, where $\odot$ denotes the element-wise multiplication, $\mathrm{attention}$ represents matrix-multiplying on the time axis by a single head of previously computed attention weights, and the linear layer recovers the dimension to the same as the input.
Bypass. The Bypass module learns channel-wise scalar weights $\mathbf{c}$ to combine the module input $\mathbf{x}$ and module output $\mathbf{y}$: $(1-\mathbf{c})\odot\mathbf{x}+\mathbf{c}\odot\mathbf{y}$. In training, we initially limit the values of $\mathbf{c}$ in range of $[0.9,1.0]$ and then change the minimum to 0.2 after 20000 steps. We found that making modules “straight-through” at the beginning (i.e. allowing very little bypass) helps model convergence..
3.3 BiasNorm
Conformer (Gulati et al., 2020) utilizes LayerNorm (Ba et al., 2016) to normalize the module activations. Given $\mathbf{x}$ with $D$ channels, LayerNorm is formulated as:
$\vspace{-0.5em}\mathrm{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mathrm{E}[\mathbf{x}]}{\sqrt{\mathrm{Var}[\mathbf{x}]+\epsilon}}\odot\bm{\gamma}+\bm{\beta}.\vspace{-0.2em}$ (1)
Specifically, it first computes the mean $\mathrm{E}[\mathbf{x}]$ and the standard-deviation $\sqrt{\mathrm{Var}[\mathbf{x}]}$ for normalizing, scaling the vector length to $\sqrt{D}$. Then it uses the learnable channel-wise scale $\bm{\gamma}$ and bias $\bm{\beta}$ for transformation, which helps to adjust the size of activations and balance the relative contributions of specific modules. However, we observe that the trained Conformer using LayerNorm suffers from two problems: 1) It sometimes sets one channel to a large constant value, e.g. 50. We argue that it aims to “defeat” the LayerNorm which fully removes the vector length, functioning as a very large value so that length information could be retained after normalization. 2) Some modules (typically feed-forward or convolution) are “dead” as they have extremely small output values, e.g., ${10^{-6}}$. We argue that early in training, the un-trained modules are not useful so they are “turned off” by the LayerNorm scale $\bm{\gamma}$ approaching zero. If the scale $\bm{\gamma}$ oscillates around zero, the inconsistent sign constantly reverses the gradient directions back-propagating to the modules. Because of the inconsistent gradient sign, the modules never learn anything useful, since this is a bad local optimum which is hard to escape because of the dynamics of stochastic gradient descent-like updates.
To address above problems, we propose the BiasNorm which is intended to be a simpler replacement of LayerNorm. Specifically, BiasNorm is formulated as:
$\mathrm{\emph{BiasNorm}}(\mathbf{x})=\frac{\mathbf{x}}{\mathrm{RMS}[\mathbf{x}-\mathbf{b}]}\cdot\exp(\gamma),$ (2)
where $\mathbf{b}$ is the learnable channel-wise bias, $\mathrm{RMS}[\mathbf{x}-\mathbf{b}]$ is the root-mean-square value taken over channels, and $\gamma$ is a scalar. We first remove the operation of mean subtraction since it is a waste of time unless it follows a non-linearity. The bias $\mathbf{b}$ serves as the large constant value which allows to retain the vector length information after normalization. Since the scale $\exp(\gamma)$ is always positive, it avoids the gradient oscillation problem.
3.4 SwooshR and SwooshL activation functions
Conformer (Gulati et al., 2020) adopts Swish (Ramachandran et al., 2017) activation function with the following formula:
$\mathrm{Swish}(x)=x\cdot(1+\exp(-x))^{-1}.$ (3)
In this work, we propose two new activation functions respectively called SwooshR and SwooshL as replacements of Swish:
$\begin{split}\mathrm{\emph{SwooshR}}(x)&=\log(1+\exp(x-1))-0.08x-0.313261687,\\ \mathrm{\emph{SwooshL}}(x)&=\log(1+\exp(x-4))-0.08x-0.035.\end{split}$ (4)
In SwooshR, the offset 0.313261687 is to make it pass through the origin; in SwooshL, the offset 0.035 was tuned, which slightly outperformed the value exactly making the curve pass through the origin. We present the curves of Swish, SwooshR, and SwooshL in Appendix Section A.2. SwooshL is roughly a right shifted version of SwooshR. Note that the suffix “L” or “R” represents whether the left or right zero-crossing is at or around $x=0$. Similar to Swish, SwooshR and SwooshL have lower bounds and are non-monotonic. Compared to Swish, the most striking difference is that SwooshR and SwooshL have non-vanishing slopes for negative inputs, which helps to escape from situations where the input is always negative and prevents the denominator term in Adam-type updates from getting dangerously small. When replacing Swish with SwooshR, we observe that the modules with bypass connections, such as feed-forward and ConvNeXt, tend to learn a large negative bias in the preceding linear layer to learn “normally-off” behavior. Therefore, we use SwooshL for these “normally-off” modules and use SwooshR for convolution modules and the rest of Conv-Embed.
3.5 ScaledAdam optimizer
We propose a parameter-scale-invariant version of Adam (Kingma & Ba, 2014) called ScaledAdam, which enables faster convergence and better performance. ScaledAdam scales each parameter’s update proportional to the scale of that parameter, and also explicitly learns the parameter scale. Algorithm 1 in Appendix Section A.1.1 presents the pseudo-code of the ScaledAdam.
Let $f(\bm{\theta})$ be the loss function that we aim to minimize, which is differentiable w.r.t. the learnable parameter $\bm{\theta}$. At each step $t$, Adam computes the parameter gradient $\mathbf{g}_{t}=\nabla_{\bm{\theta}}f(\bm{\theta}_{t-1})$, and updates the first moment $\mathbf{m}_{t}=\beta_{1}\cdot\mathbf{m}_{t-1}+(1-\beta_{1})\cdot\mathbf{g}_{t}$ and the second moment $\mathbf{v}_{t}=\beta_{2}\cdot\mathbf{v}_{t-1}+(1-\beta_{2})\cdot\mathbf{g}_{t}^{2}$ of gradients, where $\beta_{1},\beta_{2}\in[0,1)$ are coefficients used to compute the moving averages. The parameter update $\bm{\Delta}_{t}$ at step $t$ is formulated as:
$\bm{\Delta}_{t}=-\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon},\vspace{-1.5em}$ (5)
where $\alpha_{t}$ is the learning rate typically specified by an external schedule, $\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}$ is the bias-correction term, and $\epsilon=10^{-8}$. Whilst Adam is invariant to gradient scale of each parameter, we argue that it still suffers from two limitations: 1) The update $\bm{\Delta}_{t}$ in Equation 5 does not take into account the parameter scale (denoted as $r_{t-1}$). Considering the relative parameter change $\bm{\Delta}_{t}/r_{t-1}$, Adam might cause learning in relative terms too slowly for parameters with large scales, or too fast for parameters with small scales. 2) It is difficult to learn the parameter scale directly, as the direction of growing or shrinking the parameter tensor is a very specific direction in a large-dimensional space. It’s particularly difficult to shrink a parameter, since each gradient step $\mathbf{g}_{t}$ adds noise which tends to grow the parameter norm.
Scaling update. To keep the relative change $\bm{\Delta}_{t}/r_{t-1}$ over parameters of varying scales about the same, we scale the update $\bm{\Delta}_{t}$ in Equation 5 by the parameter scale $r_{t-1}$:
$\bm{\Delta}_{t}^{\prime}=-\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}.\vspace{-0.5em}$ (6)
We compute the parameter scale $r_{t-1}$ as the root-mean-square value $\mathrm{RMS}[\bm{\theta}_{t-1}]$. Because the ScaledAdam update is less prone to divergence than Adam, we use a learning rate schedule called Eden that does not have a long warm-up period; we also use absolutely larger learning rate values because the parameter RMS value is normally much less than one.
Learning parameter scale. To explicitly learn the parameter scale, we treat it as a regular parameter to be learned, as if we have factored each parameter as $\bm{\theta}=r\cdot\bm{\theta}^{\prime}$, and we are doing gradient descent on the parameter scale $r$ and the underlying parameter $\bm{\theta}^{\prime}$. Let $h$ be the gradient of the parameter scale $r$, at step $t$ we get $h_{t}=\nabla_{r}f(\bm{\theta}_{t-1})=\mathbf{g}_{t}\cdot\bm{\theta}_{t-1}^{\prime}$. Since Adam is nearly invariant to changes in the gradient scale, for simplicity we replace this with $h_{t}=\mathbf{g}_{t}\cdot(r_{t-1}\odot\bm{\theta}_{t-1}^{\prime})=\mathbf{g}_{t}\cdot\bm{\theta}_{t-1}$. Following the Adam algorithm, we maintain the first moment $n_{t}=\beta_{1}\cdot n_{t-1}+(1-\beta_{1})\cdot h_{t}$ and the second moment $w_{t}=\beta_{2}\cdot w_{t-1}+(1-\beta_{2})\cdot h_{t}^{2}$ of the scale gradients $h_{t}$. The parameter change on $\bm{\theta}$ caused by updating parameter scale from $r_{t-1}$ to $r_{t}$ is $\bm{\Delta}_{t,r}^{\prime}=(r_{t}-r_{t-1})\odot\bm{\theta}_{t-1}^{\prime}$. Similar to Equation 6, we also integrate the parameter scale $r_{t-1}$ into the update $\bm{\Delta}_{t,r}^{\prime}$:
$\begin{split}\bm{\Delta}_{t,r}^{\prime}&=-\eta\cdot\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}^{\prime}\\ &=-\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}.\end{split}\vspace{-0.99em}$ (7)
where $\eta$ is a scaling factor on learning rate $\alpha_{t}$, and we found that setting $\eta=0.1$ helps to stabilize the training. Now the update $\bm{\Delta}_{t}^{\prime}$ is replaced with $\bm{\Delta}_{t,r}^{\prime}+\bm{\Delta}_{t}^{\prime}$, which amounts to adding an extra gradient term in the direction of growing or shrinking each parameter. This also allows to simplify the network structure by removing most of normalization layers in our Zipformer Block (described in Section 3.2), since the modules now can easily learn to scale the activations in a suitable range. One similar method called weight normalization (Salimans & Kingma, 2016) decouples the parameter norm from its direction to speed up the convergence. It replaces each parameter with two parameters, respectively specifying the direction and the magnitude. However, ScaledAdam learns the parameter scales by adding an extra update term $\bm{\Delta}_{t,r}^{\prime}$, which makes writing the modeling code simpler.
Eden schedule. The proposed Eden learning rate schedule is formulated as:
$\alpha_{t}=\alpha_{\mathrm{base}}\cdot\left(\frac{t^{2}+\alpha_{\mathrm{step}}^{2}}{\alpha_{\mathrm{step}}^{2}}\right)^{-0.25}\cdot\left(\frac{e^{2}+\alpha_{\mathrm{epoch}}^{2}}{\alpha_{\mathrm{epoch}}^{2}}\right)^{-0.25}\cdot\mathrm{linear}(\alpha_{\mathrm{start}},t_{\mathrm{warmup}},t).$ (8)
Herein, $t$ is the step index, $e$ is the epoch index, $\alpha_{\mathrm{step}}$ and $\alpha_{\mathrm{epoch}}$ respectively control the number of steps and number of epochs after which we start significantly decreasing the learning rate, $\mathrm{linear}(\alpha_{\mathrm{start}},t_{\mathrm{warmup}},t)$ is a warmup scale increasing linearly from $\alpha_{\mathrm{start}}$ to 1 over $t_{\mathrm{warmup}}$ steps and then staying constant at 1, $\alpha_{\mathrm{base}}$ is the maximum value when setting $\alpha_{\mathrm{start}}=1,\alpha_{\mathrm{warmup}}=0$. The reason for making Eden dependent on both the step index $t$ and the epoch index $e$ is to keep the amount of parameter change after certain amount of training data (e.g., one hour) approximately constant when we change the batch size, so the schedule parameters should not have to be re-tuned if we change the batch size. Other versions of Eden replace the “epoch” parts of the formula with some suitable measure of the amount of data seen. In this work, we use $\alpha_{\mathrm{base}}=0.045$, $\alpha_{\mathrm{start}}=0.5$, and $t_{\mathrm{warmup}}=500$.
Efficient implementation. To speedup the optimization in ScaledAdam, we group the parameters into batches according to their shape and perform the computation batch by batch. Note that this doesn’t affect the outcome. ScaledAdam just requires a little more memory than Adam to cache the gradient moments $n_{t}$ and $w_{t}$ (in Equation 7) for the parameter scales.
4 Experiments
4.0.1 Experimental setup
Architecture variants. We build our Zipformer variants with three model scales: small (Zipformer-S), medium (Zipformer-M), and large (Zipformer-L). For the 6 encoder stacks, the numbers of attention heads are set to {4,4,4,8,4,4}, the convolution kernel sizes are set to {31,31,15,15,15,31}. In each attention head, the query dimension and value dimension are set to 32 and 12, respectively. For the three feed-forward modules in each  Zipformer block, the hidden dimensions in the first one and the last one are 3/4 and 5/4 of that in the middle one. We adjust the layers numbers, the embedding dimensions, and the hidden dimensions of the middle feed-forward in each stack to obtain different model scales:
Table 1: Configuration of Zipformer at three different scales.
Scale layer-numbers embedding-dimensions feed-forward-dimensions
S {2,2,2,2,2,2} {192,256,256,256,256,256} {512,768,768,768,768,768}
M {2,2,3,4,3,2} {192,256,384,512,384,256} {512,768,1024,1536,1024,768}
L {2,2,4,5,4,2} {192,256,512,768,512,256} {512,768,1536,2048,1536,768}
Datasets. We perform experiments to compare our Zipformer with state-of-the-other models on three open-source datasets: 1) LibriSpeech (Panayotov et al., 2015) which consists of about 1000 hours of English audiobook reading; 2) Aishell-1 (Bu et al., 2017) which contains 170 hours of Mandarin speech; 3) WenetSpeech (Zhang et al., 2022a) which consists of 10000+ hours of multi-domain Mandarin speech.
Implementation details. We use Lhotse (Żelasko et al., 2021) toolkit for speech data preparation. The model inputs are 80-dimension Mel filter-bank features extracted on 25ms frames with frame shift of 10ms. Speed perturbation (Ko et al., 2015) with factors of 0.9, 1.0, and 1.1 is used to augment the training data. SpecAugment (Park et al., 2019) is also applied during training. We use mixed precision training for our Zipformer models. We also employ the activation constraints including Balancer and Whitener to ensure training consistency and stability. The details of Balancer and Whitener are presented in Appendix Section A.3. Pruned transducer (Kuang et al., 2022), a memory-efficient version of transducer loss that prunes path with minor posterior is used as the training objective. During decoding, beam search of size 4 with the constraint of emitting at most one symbol per frame is employed (Kang et al., 2023). We don’t use external language models for rescoring, since in this work we focus on improving the encoder model. We employ word-error-rate (WER) and character error rate (CER) as evaluation metric for English and Mandarin datasets, respectively. By default, all of our models are trained on 32GB NVIDIA Tesla V100 GPUs. For Librispeech dataset, Zipformer-M and Zipformer-L are trained for 50 epochs on 4 GPUs, and Zipformer-S is trained for 50 epochs on 2 GPUs. For Aishell-1 dataset, our models are trained for 56 epochs on 2 GPUs. For WenetSpeech dataset, our models are trained for 14 epochs on 4 GPUs.
4.0.2 Comparison with State-of-the-art Models
In this section, we compare the proposed Zipformer with other state-of-the-art models.
LibriSpeech dataset. Table 2 shows the results on LibriSpeech test datasets for Zipformer and other state-of-the-art models. For Conformer, we also list the WERs reproduced by us and other open-source frameworks. Note that there is a performance gap between the open-source reproduced Conformer and the original Conformer. Our Zipformer-S model achieves lower WERs than all variants of Squeezeformer while having much fewer parameters and floating point operations (FLOPs). Our Zipformer-L outperforms Squeezeformer-L, Branchformer and our reproduced Conformer-L by a large margin while saving over 50% FLOPs. Noticeably, when trained on 8 80G NVIDIA Tesla A100 GPUs for 170 epochs, Zipformer-L achieves WERs of 2.00%/4.38% with sufficient computing resources (last row), which is the first model to approach Conformer-L to the best of our knowledge.
We also compare the speed and memory usage between the proposed Zipfomer and other state-of-the-art models. Figure 3 presents the comparison results in terms of averaged inference time and peak memory usage in inference mode for batches of 30-second audios on an NVIDIA Tesla V100 GPU. The batch size is set to 30 to ensure all models do not have out of memory problems during inference. In overall, Zipformer models achieves better trade-off between performance and efficiency than other models. Especially for the large scale, Zipformer-L requires much less computation time and memory than other counterparts.
Table 2: WER(%) comparison between different models on LibriSpeech dataset. We also include the number of parameters and FLOPs of encoder for a 30s input audio measured with DeepSpeed (Rasley et al., 2020). $*{}^{*}$rained with 8 80G NVIDIA Tesla A100 GPUs for 170 epochs.
Model Type Params (M) GFLOPs test-clean (%) test-other (%)
Squeezeformer-XS (Kim et al., 2022) CTC 9.0 18.2 3.74 9.09
Squeezeformer-S (Kim et al., 2022) CTC 18.6 33.7 3.08 7.47
Squeezeformer-SM (Kim et al., 2022) CTC 28.2 47.6 2.79 6.89
Squeezeformer-M (Kim et al., 2022) CTC 55.6 88.4 2.56 6.50
Squeezeformer-ML (Kim et al., 2022) CTC 125.1 183.3 2.61 6.05
Squeezeformer-L (Kim et al., 2022) CTC 236.3 333.7 2.47 5.97
E-Branchformer-B (Kim et al., 2023) CTC/AED 41.1 78.1 2.49 5.61
Branchformer (Peng et al., 2022) CTC/AED 116.2 238.3 2.4 5.5
E-Branchformer-L (Kim et al., 2023) CTC/AED 148.9 284.4 2.14 4.55
Conformer-S (Gulati et al., 2020) transducer 10.3 -- 2.7 6.3
Conformer-M (Gulati et al., 2020) transducer 30.7 -- 2.3 5.0
Conformer-L (Gulati et al., 2020) transducer 118.8 -- 2.1 4.3
Conformer in WeNet (Zhang et al., 2022b) CTC/AED 121.3 -- 2.66 6.53
Conformer in ESPnet (Miyazaki et al., 2023) CTC/AED 113.2 -- 2.29 5.13
Conformer-S pruned transducer 9.8 29.1 3.75 9.24
Conformer-M pruned transducer 28.4 77.0 2.96 7.11
Conformer-L pruned transducer 122.5 294.2 2.46 5.55
Zipformer-S pruned transducer 23.3 40.8 2.42 5.73
Zipformer-M pruned transducer 65.6 62.9 2.21 4.79
Zipformer-L pruned transducer 148.4 107.7 2.06 4.63
$*{}^{*}$ pruned transducer 148.4 107.7 2.00 4.38
Figure 3: (Left) Averaged inference time and (Right) peak memory usage vs. WER comparison for different models. The WER is averaged on LibriSpeech test-clean and test-other. Averaged inference time and peak memory usage are reported for the encoders in inference mode for batches of 30-second audios with batch size of 30 on a single NVIDIA Tesla V100 GPU.
Aishell-1 dataset. Table 3 shows the CERs on Aishell-1 dataset. Compared to the Conformer model implemented in ESPnet toolkit, our Zipformer-S achieves better performance with fewer parameters. Scaling up the model leads to lower WERs, and Zipformer-M/L outperform all other models.
WenetSpeech. Table 4 presents the experimental results on WenetSpeech dataset. Again, our Zipformer-M and Zipformer-L outperform all other models on Test_Net and Test_Meeting test sets. With only one third of the parameters, our Zipformer-S yields lower WERs than Conformer models.
Table 3: CER(%) comparison between different models on Aishell-1 dataset.
Model Params (M) Type Dev Test
Conformer in ESPnet (Watanabe et al., 2018) 46.2 CTC/AED 4.5 4.9
Conformer in WeNet (Yao et al., 2021) 46.3 CTC/AED -- 4.61
E-Branchformer in ESPnet (Watanabe et al., 2018) 37.9 CTC/AED 4.2 4.5
Branchformer (Peng et al., 2022) 45.4 CTC/AED 4.19 4.43
Zipformer-S 30.2 pruned transducer 4.4 4.67
Zipformer-M 73.4 pruned transducer 4.13 4.4
Zipformer-L 157.3 pruned transducer 4.03 4.28
Table 4: CER(%) comparison between different models on WenetSpeech dataset.
Model Params (M) Type Dev Test_Net Test_Meeting
Conformer in ESPnet (Watanabe et al., 2018) 116.9 CTC/AED 9.70 8.90 15.90
Conformer in WeNet (Yao et al., 2021) 116.9 CTC/AED 8.88 9.70 15.59
Conformer-MoE(16e) (You et al., 2022) 425 CTC/AED, MoE 7.67 8.28 13.96
Conformer-MoE(32e) (You et al., 2022) -- CTC/AED, MoE 7.49 7.99 13.69
Conformer-MoE(64e) (You et al., 2022) -- CTC/AED, MoE 7.19 8.36 13.72
Zipformer-S 32.3 pruned transducer 7.96 8.6 13.97
Zipformer-M 75.9 pruned transducer 7.32 7.61 12.35
Zipformer-L 160.9 pruned transducer 7.29 7.24 12.06
4.0.3 Ablation Studies
We perform ablation experiments on LibriSpeech dataset to investigate the effect of each proposed functional technique. With Zipformer-M as the base model, we make one change each time while keeping the others untouched. Table 5 presents the experimental results.
Table 5: Ablation studies for Zipformer-M, including encoder structure, block structure, normalization layer, activation function and optimizer.
Ablation Params (M) test-clean (%) test-other (%)
Zipformer-M 65.6 2.21 4.79
Encoder structure
  No temporal downsampling 94.2 2.23 5.09
Block structure
  Double Conformer-style blocks 73.9 2.18 4.95
  No NLA 58.7 2.16 4.97
  No NLA, no attention weights sharing 60.9 2.20 5.10
  No Bypass 65.5 2.25 4.86
Normalization layer
  LayerNorm 65.6 2.29 4.97
Activation function
  Only SwooshR 65.5 2.32 5.21
  Swish 65.5 2.27 5.37
Optimizer
  Adam 65.6 2.38 5.51
Encoder structure. We remove the temporal downsampling structure from Zipformer and use Conv-Embed with downsampling rate of 4 like Conformer. The resulting model has 12 Zipformer blocks with a constant embedding dimension of 512 and has more parameters than the base model. Experimental results in Table 5 show that the resulting model without the downsampled structure yields higher WERs on both test set. It indicates that the temporal downsampling structure for efficiency does not cause information loss, but facilitates the modeling capacity with less parameters.
Block structure. As each Zipformer block has roughly twice modules as a Conformer block, we replace each Zipformer block in the base model with two Conformer blocks stacked together. This leads to 0.16% absolute WER reduction on test-other even with a larger model size, suggesting the benefits of Zipformer block structure. Removing either NLA or Bypass leads to performance degradation. If we further remove the attention weights sharing mechanism after removing NLA, the model has slightly more parameters and slower inference speed, but the WERs are not improved. We hypothesize that the two attention weights inside one Zipformer block are quite consistent and sharing them does not harm the model.
Normalization layer. Replacing BiasNorm with LayerNorm in Zipformer leads to WER drops of 0.08% and 0.18% on test-clean and test-other, respectively. It indicates the advantage of the proposed BiasNorm which allows to retain some length information in normalization.
Activation function. When using only SwooshR for all modules in Zipformer, the WER drops by 0.11% and 0.42% on test-clean and test-other, respectively, which validates the effectiveness of particularly using SwooshL for the “normally-off” modules. Employing Swish leads to more performance degradation, which indicates the advantage of SwooshR over Swish.
Optimizer. When using Adam to train Zipformer, we have to apply BiasNorm for each module in Zipformer block to avoid model divergence, since Adam cannot learn the scale of each parameter to adjust the module activations like ScaledAdam. We try different learning rate factors (denoted as $\alpha_{\mathrm{base}}$) for ScaledAdam (0.025, 0.035, 0.045, 0.055) and Adam (2.5, 5.0, 7.5, 10.0) separately. Following (Gulati et al., 2020), the learning rate schedule for Adam is $\alpha_{t}=\alpha_{\mathrm{base}}\cdot 512^{-0.5}\cdot\min(t^{-0.5},t\cdot 10000^{-1.5})$. Figure A.2 in Appendix Section A.1.2 presents the averaged WERs on test-clean and test-other at different epochs as well as the learning rates at different steps. We show the best results of ScaledAdam with $\alpha_{\mathrm{base}}=0.045$ and Adam with $\alpha_{\mathrm{base}}=7.5$ in Table 5. ScaledAdam outperforms Adam by 0.17% and 0.72% on test-clean and test-other, respectively. The results indicate that ScaledAdam enables faster convergence and better performance than Adam.
5 Conclusion
In this work, we present the Zipformer, which serves as an efficient ASR encoder. It has an U-Net-like encoder structure, which downsamples the sequence to various lower frame rates. The re-designed block structure equipped with more modules reuses the computed attention weights for efficiency. It also employs the new normalization method BiasNorm, as well as the new activation functions SwooshR and SwooshL. Meanwhile, the proposed optimizer ScaledAdam enables faster convergence and better performance. Extensive experiments on LibriSpeech, Aishell-1 and WenetSpeech datasets have demonstrated the effectiveness of the proposed Zipformer.
Appendix A Appendix
A.1 ScaledAdam optimizer
A.1.1 ScaledAdam Algorithm.
Algorithm 1 ScaledAdam Algorithm. RMS refers to root-mean-square function. $g_{t}^{2}$ refers to $g_{t}\odot g_{t}$. $\alpha_{t}$ is controlled by Eden learning rate schedule. Good default settings are $\beta_{1}=0.9,\beta_{2}=0.98,\eta=0.1$, and $\epsilon=10^{-8}$.
learning rate $\alpha_{t}$; exponential decay rates for the moment estimates $\beta_{1},\beta_{2}\in[0,1)$; scaling factor on the learning rate for parameter scale $\eta$; objective function $f(\bm{\theta})$ with parameters $\bm{\theta}$; initial parameter $\bm{\theta}_{0}$.
$t\leftarrow 0$triangleright Initialize step.
$\mathbf{m}_{0}\leftarrow 0,\mathbf{v}_{0}\leftarrow 0$triangleright Initialize first and second moment of parameter gradient.
$n_{0}\leftarrow 0,w_{0}\leftarrow 0$triangleright Initialize first and second moments of parameter scale gradient.
$r_{0}\leftarrow\mathrm{RMS}(\bm{\theta}_{0})$triangleright Initialize parameter scale.
while $\bm{\theta}_{t}$ not converged do
     $t\leftarrow t+1$
     $\mathbf{g}_{t}\leftarrow\nabla_{\bm{\theta}}f_{t}(\bm{\theta}_{t-1})$triangleright Get parameter gradient.
     $h_{t}\leftarrow\mathbf{g}_{t}\cdot\bm{\theta}_{t-1}$triangleright Get parameter scale gradient.
     $r_{t-1}\leftarrow\mathrm{RMS}(\bm{\theta}_{t-1})$triangleright Update the parameter scale.
     $\mathbf{m}_{t}=\beta_{1}\cdot\mathbf{m}_{t-1}+(1-\beta_{1})\cdot\mathbf{g}_{t}$triangleright Update first moment of parameter gradient.
     $\mathbf{v}_{t}=\beta_{2}\cdot\mathbf{v}_{t-1}+(1-\beta_{2})\cdot\mathbf{g}_{t}^{2}$triangleright Update second moment of parameter gradient.
     $\bm{\Delta}_{t}^{\prime}=-\alpha_{t}\cdot r_{t-1}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}+\epsilon}$triangleright Compute parameter change.
     $n_{t}\leftarrow\beta_{1}\cdot n_{t-1}+(1-\beta_{1})\cdot h_{t}$triangleright Update first moment of parameter scale gradient.
     $w_{t}\leftarrow\beta_{2}\cdot w_{t-1}+(1-\beta_{2})\cdot h_{t}^{2}$triangleright Update second moment of parameter scale gradient.
     $\bm{\Delta}_{t,r}^{\prime}\leftarrow-\eta\cdot\alpha_{t}\cdot\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\cdot\frac{n_{t}}{\sqrt{w_{t}}+\epsilon}\odot\bm{\theta}_{t-1}$triangleright Compute parameter change by updating parameter scale.
     $\bm{\theta}_{t}\leftarrow\bm{\theta}_{t-1}+\bm{\Delta}_{t}^{\prime}+\bm{\Delta}_{t,r}^{\prime}$triangleright Update parameter.
end while
A.1.2 Comparison between ScaledAdam and Adam.
Figure A.2: Comparison between ScaledAdam and Adam in terms of: (Left) averaged WER on LibriSpeech test-clean and test-other at different epochs; (Right) learning rate at different steps.
A.2 Activation functions
Figure A.3: The activation functions: Swish, SwooshR, and SwooshL.
A.3 Activation constraints
Table 6: Ablation studies of activation constraints for Zipformer-M on LibriSpeech dataset. All models are trained for 40 epochs.
Ablation test-clean (%) test-other (%)
Zipformer-M 2.21 4.91
  No Balancer 2.23 4.97
  No Whitener 2.25 5.15
  No Balancer, No Whitener 2.25 5.07
To ensure consistency in training and avoid badly trained modules, we propose the Balancer and the Whitener, which regularize the activations in a memory-efficient way. In forward pass, they are no-ops; in backward pass, they compute the gradients of the additional losses putting constraints on the activations $\mathbf{g}^{\prime}$, and add that to the origin activation gradients $\mathbf{g}$: $\mathbf{g}=\mathbf{g}+\mathbf{g}^{\prime}$. The placements of Balancer and Whitener may not seem to follow any very clear rules. They are generally applied when encountering specific instances of not-normally-trained models or model divergence. We first locate the abnormality to a specific module and then add the Balancer or Whitener to fix it.
A.3.1 Balancer
[…]
$\begin{split}\mathcal{L}_{\mathrm{RMS}}&=|\log(\min(\max(\mathrm{RMS}[\mathbf{x}],r_{\mathrm{max}}),r_{\mathrm{min}})/\mathrm{RMS}[\mathbf{x}])|,\\ \mathcal{L}_{\mathrm{E/\sqrt{Var}}}&=|\mathrm{E}[\mathbf{x}]/\sqrt{\mathrm{Var}[\mathbf{x}]}-\mathrm{clamp}(\mathrm{E}[\mathbf{x}]/\sqrt{\mathrm{Var}[\mathbf{x}]},\mu_{\mathrm{min}},\mu_{\mathrm{max}})|,\\ \mathcal{L}_{\mathrm{balancer}}&=\mathcal{L}_{\mathrm{RMS}}+\mathcal{L}_{\mathrm{E/\sqrt{Var}}},\end{split}$ (9)
where the statistics $\mathrm{RMS}[\mathbf{x}]$, $\mathrm{E}[\mathbf{x}]$, and $\sqrt{\mathrm{Var}[\mathbf{x}]}$ are calculated in each channel. Before adding the additional gradient $\mathbf{g}^{\prime}=\nabla_{\mathbf{x}}\mathcal{L}_{\mathrm{balancer}}$ to the original activation gradient $\mathbf{g}$, $\mathbf{g}^{\prime}$ is scaled to $\mathbf{g}^{\prime}=\mathbf{g}^{\prime}\cdot\alpha/\mathrm{RMS}[\mathbf{g}^{\prime}]\cdot|\mathbf{g}|$. Herein, $\alpha$ is used to prevent $\mathbf{g}^{\prime}$ from overwhelming $\mathbf{g}$, and the per-element magnitude $|\mathbf{g}|$ is used to prevent the model from concentrating its “fixes” to the data distribution in frames with small gradients such as the padding frames. We set $\alpha=0.04$ in this work.
A.3.2 Whitener
Another failure mode on activations is that for the feature covariance, one or a few eigenvalues are dominating while others are extremely small. This tends to happen in a model that is about to diverge. Whitener encourages a more informative output distribution, by restricting the feature covariance after mean subtraction to have less unequal eigenvalues. Specifically, for output $\mathbf{x}\in\mathcal{R}^{N\times D}$ with $N$ frames of $D$-dimensional features, we first compute the covariance matrix $C=(\mathbf{x}-\mathrm{E}[\mathbf{x}])^{T}(\mathbf{x}-\mathrm{E}[\mathbf{x}])$, where $C\in\mathcal{R}^{D\times D}$, and $\mathrm{E}[\mathbf{x}]$ is per-channel mean. The auxiliary loss which measures the whiten metric $\mathcal{L}_{\mathrm{whiten}}$ is defined as:
$\begin{split}\mathcal{L}_{\mathrm{whitener}}=(\sum_{i}\lambda_{i}^{2}/D)/(\sum_{i}\lambda_{i}/D)^{2}=(\sum_{i}\sum_{j}C_{i,j}^{2}/D)/(\sum_{i}C_{i,i}/D)^{2},\\ \end{split}$ (10)
where $\bm{\lambda}=\{\lambda_{1},\dots,\lambda_{D}\}$ are the eigenvalues of the covariance matrix $C$. To keep the original activation gradient $\mathbf{g}$ dominant after adding the additional gradient $\mathbf{g}^{\prime}=\nabla_{\mathbf{x}}\mathcal{L}_{\mathrm{whitener}}$, $\mathbf{g}^{\prime}$ is scaled to $\mathbf{g}^{\prime}=\mathbf{g}^{\prime}\cdot\alpha/\ell^{2}(\mathbf{g}^{\prime})\cdot\ell^{2}(\mathbf{g})$, where $\ell^{2}$ denotes the L2 norm, and $\alpha$ is set to 0.01. The modification $\mathbf{g}=\mathbf{g}+\mathbf{g}^{\prime}$ is done only when the whiten metric $\mathcal{L}_{\mathrm{whiten}}$ is above a certain value $w_{\min}$ to prevent the model from learning pathological activation distributions. We usually set $w_{\min}$ to 10.
A.3.3 Ablation studies.
We perform ablation experiments on LibriSpeech dataset to validate the effect of Balancer and Whitener. Table 6 presents the experimental results. All models are trained for 40 epochs. Removing Balancer does not lead to obvious change on model performance. However, it would increase the risk of model divergence without the value range constraints especially when employing mixed precision training. Removing Whitener results in 0.04% and 0.24% WER reduction on test-clean and test-other, respectively. This indicates that restricting the feature covariance to have less unequal eigenvalues in Whitener can boost performance.
A.4 Experiments on LibriSpeech dataset
A.4.1 Training configurations of Zipformer models
Before training, the Mel filter-bank features are per-computed and saved to disk. In training, we use DynamicBucketingSampler in Lhotse toolkit (Żelasko et al., 2021) to form the batches, where the batch size is determined dynamically given the constraint of the maximum total speech duration (in seconds). Table 7 presents the training configurations of Zipformer models on LibriSpeech dataset with speed perturbation with factors of 0.9, 1.0, and 1.1.
Table 7: Training configurations of Zipformer models on LibriSpeech dataset.
Model Type Params (M) Max duration (s) GPUs Epochs Training time / epoch (m)
Zipformer-S CTC 22.1 1700 2 32G Tesla V100 100 86
Zipformer-M CTC 64.3 1400 4 32G Tesla V100 100 60
Zipformer-L CTC 147.0 1200 4 32G Tesla V100 100 76
Zipformer-S CTC/AED 46.3 1700 2 32G Tesla V100 50 105
Zipformer-M CTC/AED 90.0 1200 4 32G Tesla V100 50 67
Zipformer-L CTC/AED 174.3 1200 4 32G Tesla V100 50 84
Zipformer-S pruned transducer 23.3 1500 2 32G Tesla V100 50 87
Zipformer-M pruned transducer 65.6 1000 4 32G Tesla V100 50 69
Zipformer-L pruned transducer 148.4 1000 4 32G Tesla V100 50 80
Zipformer-L pruned transducer 148.4 2200 8 80G Tesla A100 200 18
A.4.2 Comparison with state-of-the-art models
As an extension of Table 2, Table 8 adds the results on LibriSpeech dataset for Zipformer with CTC and CTC/AED architectures respectively. For the Zipformer CTC/AED model, we use a 6-layer Transformer as AED decoder, each layer with attention dimension of 512, attention heads number of 8, and feed-forward hidden dimension of 2048. The Zipformer CTC models are trained for 100 epochs while the Zipformer CTC/AED models are trained for 50 epochs. Detailed training configurations are provided in Section A.4.1.
For the CTC systems, Zipformer-M outperforms Squeezeformer-ML on both test sets with only about half the number of parameters, and Zipformer-L also surpasses Squeezeformer-L by 0.27% on test-other with fewer parameters. For CTC/AED systems, Zipformer-M outperforms Conformer models and Branchformer, while Zipformer-L achieves comparable results with E-Branchformer-L. Note that as presented in Figure 3, Zipformer-L is much more efficient than E-Branchformer-L.
Table 8: WER(%) comparison between different models on LibriSpeech dataset. We also include the number of parameters and FLOPs of encoder for a 30s input audio measured with DeepSpeed (Rasley et al., 2020). $*{}^{*}$rained with 8 80G NVIDIA Tesla A100 GPUs for 170 epochs.
Model Type Params (M) GFLOPs test-clean (%) test-other (%)
Squeezeformer-XS (Kim et al., 2022) CTC 9.0 18.2 3.74 9.09
Squeezeformer-S (Kim et al., 2022) CTC 18.6 33.7 3.08 7.47
Squeezeformer-SM (Kim et al., 2022) CTC 28.2 47.6 2.79 6.89
Squeezeformer-M (Kim et al., 2022) CTC 55.6 88.4 2.56 6.50
Squeezeformer-ML (Kim et al., 2022) CTC 125.1 183.3 2.61 6.05
Squeezeformer-L (Kim et al., 2022) CTC 236.3 333.7 2.47 5.97
E-Branchformer-B (Kim et al., 2023) CTC/AED 41.1 78.1 2.49 5.61
Branchformer (Peng et al., 2022) CTC/AED 116.2 238.3 2.4 5.5
E-Branchformer-L (Kim et al., 2023) CTC/AED 148.9 284.4 2.14 4.55
Conformer-S (Gulati et al., 2020) transducer 10.3 -- 2.7 6.3
Conformer-M (Gulati et al., 2020) transducer 30.7 -- 2.3 5.0
Conformer-L (Gulati et al., 2020) transducer 118.8 -- 2.1 4.3
Conformer in WeNet (Zhang et al., 2022b) CTC/AED 121.3 -- 2.66 6.53
Conformer in ESPnet (Miyazaki et al., 2023) CTC/AED 113.2 -- 2.29 5.13
Conformer-S pruned transducer 9.8 29.1 3.75 9.24
Conformer-M pruned transducer 28.4 77.0 2.96 7.11
Conformer-L pruned transducer 122.5 294.2 2.46 5.55
Zipformer-S CTC 22.1 40.8 2.85 6.91
Zipformer-M CTC 64.3 62.9 2.51 6.02
Zipformer-L CTC 147.0 107.7 2.49 5.7
Zipformer-S CTC/AED 46.3 40.8 2.46 6.04
Zipformer-M CTC/AED 90.0 62.9 2.22 4.97
Zipformer-L CTC/AED 174.3 107.7 2.09 4.59
Zipformer-S pruned transducer 23.3 40.8 2.42 5.73
Zipformer-M pruned transducer 65.6 62.9 2.21 4.79
Zipformer-L pruned transducer 148.4 107.7 2.06 4.63
$*{}^{*}$ pruned transducer 148.4 107.7 2.00 4.38